Shift report 20011111 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Nov 11 2001 - 23:51:47 EST

  • Next message: Apache: "Shift report 20011112 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Nov 11 23:51:46 2001, Shift: 16:00-24:00
    Supervisor: BHS, with: 
    Log book#10 pages 127 - 130
    
    Shift Summary:
    ------------------------------------------------------------
    1: We got a good store at 17:35, which lasted until 23:30. 
    PHENIX and BRAHMS voted to dump around 22:00, but noone
    listens to us...
    
    2:  We lost the first, really good part of the store because
    rack 2 in the FE-hut went dead. We investigated and found
    that the fuse on the back power connector panel had gone,
    and reset it. The problem reappeared several times, and
    in the end we found that some of the power connectors had
    died. We moved the cables to new, working connectors and in
    the end the thing stabilized. We did not find out what caused
    it to trip/short in the first place.
    
    So: Runs 5763 and 5764 are data runs, but contain no H1 and
    variable HV on the other detectors. They should probably
    just be disregarded.
    Proper runs:
             Trig 4    Trig 6    Settings
    5765     193k      102k      MRS@90, D5@4A
                                 FFS@30, D1-D4@1/15B
    5766     298k       91k      Same
    5767      88k       27k      MRS@40, D5@4A, FS unchanged.
                       ----
                       221k
    
    3: No problems apart from what was mentioned above. MRC did 
    call once and inform us that the store that was lost at the
    end of last run probably died because of our magnet problems.
    It hit the wall as the engineer switched on D2, and for some
    reason they think this interfered with DX2...
    
    That's about it :-)
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Nov 11 2001 - 23:52:41 EST